The stock had returns of -50% and +100% in two consecutive years.

If the initial stock was $100, what is it at the end of the first and second year?

In the first year you invest in 1 share, in the second year in another 2 shares

Find:

  1. Arithmetic Average
  2. Geometric Average
  3. Dollar Weighted Average

Which of the averages best reflects historical performance over long time?

Given above data, what return would be reasonably expected for the third year?
Answer #1

Stock Value = $100

1st year return = -50% of $100 = $50

2nd year return = 100% of $50 = $100

Now,

Arithmatic Avg = (150 + 50 + 50)/3 = $83.333

Geometric Avg = squareroot ( 150*100) = $122.5

Dollar weighted Avg best reflects historical performance over log time.

for third year

more than 100% return expected for the third year.